Montserrat Volcano Observatory, Montserrat, West Indies

Scientific Report 55
01 March 1997

This report covers the 7-day period to 24:00 on 01 March 1997


Earthquake swarms have continued to dominate the seismicity but at a slightly reduced level to that seen last week. The dome complex has continued to grow with a swap to vertical growth of the January 20 dome. The eastern face has continued to be unstable producing rockfalls and small pyroclastic flows.

Visual Observations

The viewing conditions have been generally poor this week, but brief views of the crater area have been possible. At the start of the period, few changes in the dome were visible. Major vertical growth of the January 20 dome was seen for the first time on 27 January. There have been few pyroclastic flows.

On 23 February, new growth was seen at the top of the eastern face of the dome. Two large, unstable blocks were noted, with one of the blocks being about 30 metres high. By 27 February, the next time that the dome was visible, these blocks had disappeared.

Deep gullies observed on the eastern face in the previous week had become less pronounced, and on 27 February it was clear that the main growth in the previous few days had been vertical, rather than the outward growth of the eastern face which has characterised the last few weeks. A small spine was noted at the top of the dome. Vertical growth of the dome was about 42 m between 18 February and 1 March, when a dome survey was undertaken.

The north side of the dome became more active during the week, with more small pyroclastic flows observed from there, creating a small scar.

The slow crumbling of Galway's Wall continued, with the development of a fourth chute in the top of the wall, to the east of the other chutes. Small rockfalls from the dome continued to fall over the wall, although the maximum run-out distance has not increased for over two weeks.


The level of seismic activity has been moderate, with a continuation of regular swarms of earthquakes at a slightly reduced rate compared with the previous week, and a very slight increase in the number of rockfall signals.

The level of high-frequency tremor, recorded only at the Gages seismometer, has been high at times. This tremor, previously referred to as broadband tremor, has frequencies above 3 Hz, has been observed throughout the eruption and is thought to be due to a local source. No clear correlation has ever been observed between this tremor and visual phenomena.

Three short episodes of continuous tremor were recorded, each lasting for 1 to 2 hours, on 26 to 28 February. The episode on 28 February was unusual in that it did not follow on from an earthquake swarm.

Table 1: Earthquake types
These earthquake counts are of events that triggered the short-period seismic network event recording system between 0000 and 2400 each day.

Date	VT	LP	Hybrid	Dome RF
23 FEB 97	4	0  	75  	18
24 FEB 97	39	0  	30  	12
25 FEB 97	24	0	4  	14
26 FEB 97	39	0  	40  	15
27 FEB 97	28	0  	58  	33
28 FEB 97	6	0  	15  	25
01 MAR 97	101	0  	62  	6

Table 2: VT / hybrid earthquake swarms and tremor episodes

Date	    Start time	Durn (hrs)	VTs	Hybrids	Gap* (hrs)
22 FEB 97 	19:00 	9.80  	41	83	47.0
24 FEB 97 	17:57 	8.33  	49	30	39.0  
26 FEB 97 	09:10 	7.55  	38	37	20.0
27 FEB 97 	05:06 	7.97  	29	61	26.0
28 FEB 97 	07:12 	5.07  	6	14	19.0
01 MAR 97	01:56 	17.10  	98	58	20.0
01 MAR 97 	21:54 	17.95  	21	95	
* Gap is time interval between start of swarms

Ground Deformation

The extensometer which has been in place since mid-December failed on 24 February, after being covered by a thick ash deposit. Crack 2 on Chances Peak was measured on the 1st of March. It had opened by a further 2cm since the last measurement on the 11th of February and it has opened by a total of 10cm throughout the measurement period. The dextral shear movement on this crack had increased by 8cm. The total shear displacement is now 16cm which has essentially all occurred since 28th of January.

A GPS survey of the volcano's EASTNET (stations at Harris, Whites, Tar River, Windy Hill and Farrells) was carried out on 26th of February. Baseline lengths and station heights were all within confidence limits of the long term means.

Dome Volume Measurements

A dome survey was carried out on 1 March which concentrated on the January dome and the north face where minor activity was still occurring. From the DEM's produced, growth since the 18 February has again been concentrated within with the January scar although this time extrusion has mostly occurred at the top of the dome. The height of the dome has increased by 38m and the talus slope directly to the east of this high point has increased in thickness by 40m. As there has been very low level of rockfall activity during this period and as the talus/solid rock contact is at a lower height than during the last survey it is thought that the build up of material in this area is due to the solid lava spreading out to the east as opposed to just a build up in the talus material. About 10m of uplift can be seen in a wedge of the October dome directly adjacent to the top of the January bulge on the northern side of the contact. This wedge appears to have been thrust up by the neighbouring January extrusion.

In the north east in the area seen to have been shedding material, 20m of loss can be detected and in the far north a small patch of uplift of about ten metres can be seen. Although this face remains relatively quiet small changes have been consistently detected throughout the last few surveys and minor endogenous growth is thought to be occurring here.

The volume increase obtained from this survey since 18 February was 3.42 million cubic metres. For a total volume this has been added to the 40.2 million obtained for the estimate of 17 February (this is the lower of the two estimates calculated for that date as it is considered to be the more consistent with previous surveys). The new volume now stands at 43.7million cubic metres and the average extrusion rate for that 11 day period is 4.37m3/s. However visual observations suggest that the bulk of the new growth at the top of the January dome occurred within the period 22 -28 February, if one assumes baseline extrusion rate for the remainder of the survey period , extrusion rates for 22-28 February could have been as high as 5.8m3/s. Dome profiles from Whites show an increased progradation of the talus slope to the east and that the lateral spreading of the dome established in the last few weeks is continuing but that the January dome has a significant bulge on the top. Active extrusion appears to de occurring from the back of the January scar thrusting large curved slabs up and to the east.

Gas Measurements

The Mini-COSPEC machine was returned on 24 February and measurements of sulphur dioxide flux were made on 25 February (217 t/d) and 1 March (165 t/d). These values are lower than the last values recorded at the beginning of February (300 to 700 t/d)..

Sulphur dioxide diffusion tubes were collected on 23 February and sent to the UK for analysis on 24 February. A further set of tubes were placed at 4 locations to the west of the volcano and near Whites to the east of the volcano.

Environmental Monitoring

Rain water samples were collected at 3 locations around the volcano on 23 February, and results are in Table 3. One sample was also collected from run-off water by the side of the road to Upper Amersham and one sample was collected from the overflow to a water storage tank at Fairfield.

These results show that the rainwater directly west of the volcano continues to be highly acidic and has high concentrations of certain anions. The water from the storage tank has anion concentrations and pH levels well within WHO guidelines for drinking water.

Table 3: Rain and surface water geochemistry, 23 Feb 97
Units: conductivity (mS/cm), total dissolved solids (g/l), sulphate (mg/l), chloride (mg/l), fluoride (mg/l),

Location		pH	Cond.	TDS	Slphts	Chlrds	Fluorides
Upper Amersham		2.60	1.134	0.567	16	166	>1.5
Lower Amersham		2.78	0.756	0.378	17	420	>1.5
Police HQ, Plymouth	2.93	0.463	0.231	nd	45	>1.5
Amersham Stream		3.99	2.85	1.42	nd	500	1.3
Fairfield tank		7.74	0.604	0.302	39	57	0.4

No results on ash analyses were received during this period.

Staff Changes
Rod Stewart, BGS

Kenmuel Spence, Soufriere Monitoring Unit, St Vincent and the Grenadines

Montserrat Volcano Observatory